OpenStudy (anonymous):

solve the equation. 3x^-2 - 14x^-1 - 5 = 0

OpenStudy (anonymous):

3x^-2 - 14x^-1 - 5 = 0 let x^-2=v^2,x^-1=v 3v^2 - 14v - 5 = 0, solving this quadratic equation v=5, -1/3 now subs x=1/v =1/5 ans..... and x=1/v=-3 ans.....